What is the wavelength of an earthquake that shakes you with a frequency of 29 Hz and gets to another city 69 km away in 24.43 s?


Round your answer to 0 decimal places


1
Expert's answer
2021-04-26T17:03:59-0400
"\lambda=\frac{v}{f}=\frac{s}{ft}\\\lambda=\frac{69000}{29\cdot24.43}=97\ m"